OpenStudy (anonymous):

Three semi circles are drawn outside a triangle so that the sides of the triangle are the diameters of the semi circles. If the triangle is a right angled triangle and the semi circles have areas 18pie, 32pie and 50pie, determine the are of the triangle.

OpenStudy (king):

48?

OpenStudy (anonymous):

how did you get that?

OpenStudy (king):

wait ill explain

OpenStudy (anonymous):

thanks

OpenStudy (king):

See,we know that area of one circle is pi r^2 yes or no

OpenStudy (king):

yes or no?

OpenStudy (anonymous):

yes

OpenStudy (king):

So, diameters of the 3 circles is the same as the 3 sides and pi r^2 can be written as (d/2)^2 yes or no?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you do the equation? please

OpenStudy (king):

so calculate the diameter for all 3 circles I shall work it out for one circle \[18\pi=\pi \times d ^{2}/4\] so pi is cancelled \[18 \times 4=d ^{2}\] \[d=\sqrt{72}\] that is 6sqrt2

OpenStudy (king):

work out for remaining 2 and tell me what u get be sure to get it as something sqrt 2

OpenStudy (king):

are u doing it ?

OpenStudy (king):

anyways after finding all 3 sides or diameters find semi-perimeter semi perimeter is perimeter by 2 then use heron's formula to find area heron's formula is \[\sqrt{s(s-a)(s-b)(s-c)}\] where s is semi-perimeter

OpenStudy (king):

ok did u understand do u want me to explain some part?

OpenStudy (anonymous):

is there another formula I can use?

OpenStudy (king):

none that i know of sry

OpenStudy (anonymous):

ok

OpenStudy (king):

ok so u understand everything rite?

OpenStudy (anonymous):

yes but how do I get 48?

OpenStudy (king):

find all 3 sides they should be \[6\sqrt{2},8\sqrt{2},10\sqrt{2}\] s=\[24\sqrt2/2=12\sqrt2\] area=\[\sqrt{12\sqrt2\times 12\sqrt2-6\sqrt2 \times 12\sqrt2- 8\sqrt2 \times 12\sqrt2-10\sqrt2}\] that is \[\sqrt{12\sqrt2 \times 6\sqrt2 \times 4\sqrt2 \times 2\sqrt2}\] that is sqrt2 squared *24squared that is 24*2=48

OpenStudy (anonymous):

thanks a lot ur the best also my hero